Circular No. 7152

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 H3 (LINEAR) Additional observations are available on MPEC 1999-H29, from whence

the following preliminary orbital elements are taken:

T = 1999 Aug. 17.907 TT  ω = 101.^{o}801
Ω = 332.669 }2000.0
q = 3.50186 AU  i = 115.832
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Apr. 22 19^{h}17.^{m}68 +44^{o}02.^{′}9 3.548 3.665 88.^{o}6 15.^{o}9 15.4
27 19 11.41 +45 54.0 3.491 3.651 91.1 16.0 15.3
May 2 19 03.83 +47 44.4 3.439 3.639 93.3 16.0 15.3
7 18 54.81 +49 32.5 3.392 3.627 95.3 16.1 15.2
12 18 44.22 +51 16.3 3.351 3.615 96.9 16.1 15.2
17 18 31.94 +52 53.5 3.317 3.604 98.3 16.1 15.2
22 18 17.96 +54 21.5 3.290 3.594 99.2 16.1 15.1
27 18 02.30 +55 37.9 3.270 3.584 99.8 16.2 15.1
June 1 17 45.15 +56 40.3 3.257 3.574 99.9 16.2 15.1
6 17 26.79 +57 26.4 3.251 3.565 99.7 16.3 15.1
11 17 07.66 +57 55.0 3.253 3.557 99.0 16.4 15.1

SUPERNOVA 1999bw IN NGC 3198 A. V. Filippenko, W. D. Li, and M. Modjaz, University of California

at Berkeley, write: “A CCD spectrum (range 450–710 nm, resolution 0.7 nm) of the reported possible supernova in NGC 3198 (cf. IAUC 7149, 7150), obtained on Apr. 24 with the 3-m Shane reflector at Lick Observatory looks very similar to that of the type-IIn SN 1997bs in NGC 3627 (cf. IAUC 6627). The narrow Balmer emission lines (FWHM ∼ 600 km/s) have a broader base with FWZI roughly 3000 km/s, and there is a hint of weak Fe II emission near 520 nm. Like SN 1999bw, SN 1997bs was quite subluminous (M ∼ −13). Another example of an object having a similar spectrum is SN 1999Z (cf. IAUC 7107), but it was much more luminous. The type-IIn subclass clearly spans a very broad range of properties, as was already known to some extent (e.g., Filippenko 1997, Ann. Rev. Astron. Astrophys. 35, 309). It remains to be seen whether the explosion mechanism is core collapse in all of these objects; perhaps in some cases the star is not fully destroyed (e.g., SN 1961V; Filippenko et al. 1995, A.J. 110, 2261).”

1999 April 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green